log3(5x+5) - log3(x2 - 1) = 0

Answer by mukhopadhyay(490)

log3(5x+5) - log3(x2 - 1) = 0
=>log(base3) [5(x+1)]/(x^2-1) = 0
=> 5(x+1)/[(x+1)(x-1)] = 1
=> 5/(x-1) = 1
=> x=6
Answer: x=6